OpenStudy (anonymous):

A gym offers regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? i get the problem i just can't understand how I'm supposed to know who was a regular member and who was a off peak member

OpenStudy (anonymous):

@KlOwNlOvE

OpenStudy (asevilla5):

Call the $80 membership x, and the $60 membership y. How do you express the folloing line as an equation using x and y? "the gym sold a total of 420 memberships "

OpenStudy (anonymous):

\[80x+60y=420\]

OpenStudy (anonymous):

solve for x so substitute 0 in for 60 and divide 420 by 80

OpenStudy (asevilla5):

x is the number of $80 memberships sold y is the number of $60 memberships sold x + y represents the total number of memberships sold at both prices, right? x + y is the total number of memberships sold. We are told that a total of 420 memberships were sold, so that gives us one equation. x + y = 420 For the second equation, we deal with the amount the memberships cost and the total amount of money received. Since x represents the number of $80 memberships sold, the amount of money that was received by selling $80 memberships is 80x. The same is true of the $60 memberships. y number of them were sold at $60 each, so the amount of money received from selling $60 memberships is 60y. We are told the total amount of money received was $31,100. This is the info we need to find the second equation. 80x + 60y = 31100 Now we have two equations that need to be solved simultaneously. The system of equations is: x + y = 420 80x + 60y = 31100

OpenStudy (asevilla5):

Now we need to solve the system of equations. Let's divide the entire second equation by 10. x + y = 420 8x + 6y = 3110 Now let's multiply the first equation by -6 -6x - 6y = -2520 8x + 6y = 3110 Now we add the equations. 2x = 590 x = 295 Since x represents the number of $80 memberships, the answer is 295

OpenStudy (anonymous):

^ got the last number mixed up

OpenStudy (anonymous):

ohhhh ok

OpenStudy (anonymous):

thank you both!